OpenStudy (anonymous):

Find the equation of the line that contains the point whose coordinates are (10, -1) and has slope 4/5.

hero (hero):

y - y1 = m(x-x1) y +1 = 4/5(x - 10) y = 4/5x - 8-1 y = 4/5x - 9

hero (hero):

Any questions?

hero (hero):

Hi